OpenStudy (anonymous):

Find the next three terms of the sequence. Then write a rule for the sequence. 648, 216, 72, 24 -119 -547 -1449 given a number of finite in term A sequence infinfely sequence

OpenStudy (anonymous):

notice that each number is one-third of its previous number. This is a ______ series.

OpenStudy (anonymous):

I saw you post this question 3 times and got the instruction above. It seems you still not know how to step up? 1st number is 648 2nd number is 648/3 = 216 3rd number is 216/3=72 4th number is 72/3= 24 5th number is 24/3 = 8 6th number is 8/3 = 2.67 7th number is 2.67 /3 =0.89 and the general form of this is a_n = a_n-1/3 or if you don't like that form, you can use recursive to construct the form as a_n = 648* (1/3)^n . if you still not get. don't worry, just make question, I'll be back later

OpenStudy (anonymous):

hi friend, I have no idea about yours, I just follow the instruction above. I "translate" it to you . I saw after getting that instruction, you still make the same question in a new post. So I guess so.

OpenStudy (anonymous):

Suggestion: if you don't get, make more question, that's it. It seems you are a so shy girl. disappear again when I am talking to you. I'm ok with that, girl!

OpenStudy (anonymous):

thank you but i was busy early

OpenStudy (anonymous):

ok, let it go.